### v1

# FutureGen: LLM-RAG Approach to Generate the Future Work of Scientific Article

Abstract
The future work section of a scientific article outlines potential research directions by identifying gaps and limitations of a current study. This section serves as a valuable resource for early-career researchers seeking unexplored areas and experienced researchers looking for new projects or collaborations. In this study, we generate future work suggestions from key sections of a scientific article alongside related papers and analyze how the trends have evolved. We experimented with various Large Language Models (LLMs) and integrated Retrieval-Augmented Generation (RAG) to enhance the generation process. We incorporate a LLM feedback mechanism to improve the quality of the generated content and propose an LLM-as-a-judge approach for evaluation. Our results demonstrated that the RAG-based approach with LLM feedback outperforms other methods evaluated through qualitative and quantitative metrics. Moreover, we conduct a human evaluation to assess the LLM as an extractor and judge. The code and dataset for this project are here: code : HuggingFace
1 Introduction
The future work section in a scientific article plays a crucial role in amplifying a study’s impact by demonstrating foresight and highlighting the broader implications of research Nicholas et al. (2015); Aguinis et al. (2018). It serves as a catalyst for further exploration, interdisciplinary collaboration, and new ideas, transforming a single study into a foundation for future advancements Al Azhar et al. (2021); Hara et al. (2003). Beyond academia, future work insights benefit policymakers and funding agencies by identifying emerging research directions and prioritizing areas for strategic resource allocation Hyder et al. (2011); Thelwall et al. (2023); Simsek et al. (2024). A well-constructed future work section serves both methodological and practical purposes: it encourages researchers to critically reflect on their study’s limitations, fostering higher-quality subsequent research while also streamlining the peer review process by clarifying the authors’ awareness of challenges and next steps Kelly et al. (2014). Acknowledging a study’s limitations provides clear signals for further exploration, bridging current findings with future advancements Conaway et al. (2015). Additionally, analyzing the evolution of future work trends helps researchers align with current priorities, uncover unexplored gaps, and avoid redundancy. Furthermore, understanding long-term research trajectories allows early-career researchers to identify high-impact topics and strategically position their contributions in advancing scientific progress Ortagus et al. (2020).
Author-written future work sections have many problems. Firstly, they are often unspecific, ambiguous, not always easy to find, and speculative in nature Suray et al. (2024). Secondly, authors may struggle to articulate meaningful future directions, particularly when faced with space constraints or time limitations. Thirdly, they may hesitate to share research plans without a clear reward. This may partially explain why many future work proposals are overlooked post-publication Teufel (2017).
Advances in artificial intelligence (AI) offer transformative potential for addressing this gap. Unlike traditional methods, AI can systematically synthesize research trajectories, uncover latent connections, and propose novel directions that align with emerging trends Wang et al. (2023). For example, Si et al. (2024) shows that LLM-generated ideas are more novel than human expert-generated ones. However, current applications of AI like ChatGPT risk homogenizing outputs and reducing individual creativity Ashkinaze et al. (2024); Anderson et al. (2024). Standard LLMs may generate overgeneralized, irrelevant, or fabricated future work directions. To address these challenges, this work utilizes an LLM to suggest future work, enhances its output using LLM-based feedback, and incorporates RAG using cross-domain insights.
Evaluation of AI-generated future work sentences is challenging. Traditional Natural Language Processing (NLP) evaluation metrics that rely on n-gram text overlaps or semantic similarity, such as ROUGE Lin (2004), BLEU Papineni et al. (2002), and BERTScore Zhang et al. (2019) to compare the generated text with references, fail to fully capture the nuances of this generation process. To address this issue, we incorporate LLM-based evaluation providing human-like assessments with explanations. If the generated future work does not meet quality thresholds, we refine it using an LLM-driven feedback loop, improving its alignment with the context. We mitigate issues like vagueness and redundancy observed in prior AI-driven ideation tools by integrating iterative LLM feedback. Beyond generation, our approach embeds temporal trend analysis to track evolving shifts over time. Our Contributions can be summarized as follows:
  • Dataset. We create a dataset of future work sentences from nearly 8,000 papers collected from the ACL conference from 2012 to 2024. Papers often do not have exclusive future work sections (they are combined with general conclusions or limitations), so we use LLM to extract appropriate sentences. We then validated this extraction process with human annotators, ensuring that the LLM accurately identified future-work content.
  • Trend Analysis. We analyze how future work has evolved in the NLP domain over the past 13 years by generating topics and titles using LLM-based Topic modeling. Traditional approaches that rely only on topic words or clustering do not provide a nuanced, interpretable picture. By integrating LLM-based topic modeling with BERTopic Grootendorst (2022), our method generates not only topics but also descriptive titles, offering a more meaningful and comprehensive analysis of the trends in future work.
  • Future Work Generation. We use LLMs to generate future work sentences for NLP papers. We overcome the context window length problem by selecting key sections. Our RAG system incorporates additional relevant information from other papers, enhancing the quality and depth of the generated future work
  • LLM as a Judge and Feedback Mechanism. We incorporate an LLM as a judge to evaluate the quality of the generated text rather than only relying on NLP evaluation metrics and use the judge feedback to improve the generation process.
Figure 1: Overview of the future work Generation System using LLM and RAG.
2 Related Work
Recent advancements in NLP and large language models (LLMs) have enabled the automatic extraction and generation of various sections of scientific articles, such as Abstracts Gonçalves et al. (2018), Methodologies Houngbo and Mercer (2012), and Limitations Al Azher et al. (2024); Azher (2024); Al Azher and Alhoori (2024). In the domain of future work, prior studies have focused on tasks such as extraction Hu and Wan (2015), classification Zhang et al. (2023), identifying creative topics Song et al. (2021), thematic categorization Hao et al. (2020), and trend prediction Qian et al. (2021). For example, a BERT-based model has been used to annotate “future research” sentences, enabling the extraction, clustering Zhu et al. (2019), and impact analysis of these sentences Suray et al. (2024). Other work has integrated RAG for idea generation Radensky et al. (2024) and further trend prediction Qian et al. (2021). However, most efforts have concentrated on identifying, collecting, or forecasting trends rather than on synthesizing suggestive future work that provides actionable insights for researchers. Our work addresses this gap by leveraging LLMs to generate suggestive future work from key sections of research papers. By incorporating RAG-augmented cross-paper insights, our approach enhances the relevance and coherence of the generated suggestions.
The application of LLMs in scientific discovery has gained significant attention, particularly in generating novel research ideas and hypotheses. For example, new scientific discoveries in the biomedical field Qi et al. (2023), LLM-based agents are used to automatically generate and test social scientific hypotheses Manning et al. (2024), and a probabilistic model is used for hypothesis generation and scientific exploration using reward functions Jain et al. (2023). Moreover, LLMs are capable of generating novel research ideas Lu et al. (2024). Statistical tests show that LLM-generated ideas exhibit greater novelty compared to human-generated ones after extracting research topics from recent conferences and prompting both LLMs and humans to generate ideas Si et al. (2024). While prior studies have explored LLM-driven hypothesis generation in other fields, such as biomedical, automated suggestive scientific exploration and idea generation in the NLP domain, remain undiscovered.
A critical aspect of improving LLM-generated content is the use of human feedback mechanisms Ouyang et al. (2022). In its absence, self-refinement techniques have been proposed, such as a self-debugging framework where LLMs identify and correct their mistakes without human intervention Chen et al. (2023). Moreover, a model that generates instructions, inputs, outputs, and filters invalid results with fine-tuning itself iteratively Wang et al. (2022). AI feedback has been used for in-context learning with prompt criticism Fu et al. (2023), in a GPT-4-based three-feedback-mechanism system to generate novel hypotheses Yang et al. (2023), and in LLM-generated responses with utility-based feedback for prompt revision Peng et al. (2023).
Our work builds on these advancements by using LLMs not only to generate future work sections using LLM but also to act as a judge for scoring and justification of the output. Inspired by self-refinement techniques Madaan et al. (2023), we iteratively incorporate LLM feedback to refine the input, eliminating the need for additional training or reinforcement learning and thereby improving the quality of the generated content.
Evaluating LLM-generated text is vital for scientific writing. Studies comparing human and LLM performance Chiang and Lee (2023); Nguyen et al. (2024) show that LLM-based evaluations (covering grammaticality, cohesiveness, accuracy, coverage, etc.) align well with human judgments. Building on these insights, we use LLM-based evaluation to assess the quality of generated future work. In summary, while some progress has been made in automating the generation of scientific article sections and leveraging LLMs for scientific discovery, the suggestive generation of future work sections remains an open challenge. Our work addresses this gap by proposing a novel approach that combines LLM generation with RAG, self-refinement, and robust evaluation metrics.
3 Dataset Collection and Extraction
We extracted future work sentences from approximately $8,000$ papers published in the ACL conference from 2012 to 2024 using the Science Parse [^1] https://github.com/allenai/science-parse tool. The extraction process is as follows: (1) If a paper has a section explicitly labeled “Limitations and future work”, we extract the entire section. (2) If “future work” is not in the section title, we extract sentences with at least one of the phrases “future” or “future work”. Such regex-based string matching has a high recall, and we further filter these sentences using LLMs to improve precision (Details  5.2) (§4). This produces future work paragraphs from 6,227 papers, averaging five sentences per paper with an average word length of 65.
4 Methodology
Our methodology consists of two steps:
(1) Generating Topics and Titles of Future Work: We applied Topic Modeling with LLM to generate topics and titles for trend analysis of future work Al Azher et al. (2024).
(2) Generating future work: We used an LLM-based RAG approach to generate future work from the key sections of research papers with LLM-based evaluation and feedback.
4.1 Generating Topics and Titles of future work
Our dataset spans 13 years, with each year containing $n$ research papers. We extracted its future work section, denoted as $FW={fw_{1},fw_{2},....fw_{n}}$, where $fw_{i}$ represents the ‘future work’ paragraph containing a bunch of sentences from paper $p_{i}$. We then input these texts into BERTopic, which clusters them into $x$ topics (e.g., $t_{1},t_{2}...t_{x}$). Each topic $t_{i}$ comprises a set of topic words $tw_{1},tw_{2}...tw_{z}$ and a corresponding collection of relevant texts $R_{i}$. We utilized a topic modeling approach named BERTopic, utilized a transformer-based embedding from BERT convert to dense vector, and then created clusters to identify groups of similar documents. To enhance interpretability, we integrate an LLM) with BERTopic to generate a descriptive title for each topic (see Figure 3 in the Appendix). This entire process is repeated for every year in our dataset and collected topic title for each year (Table 9, 7, Appendix).
4.2 Generating Future Work
Figure 1 illustrates the workflow of our proposed model. We choose $X$ random papers from the dataset and store them into a vector database. Later, we select the top three sections most similar to "future work" using cosine similarity (Task 1) (step 1a) and send them to the Retriever, which gathers relevant information from a vector database (steps 2a-5). The augmented input is then processed by the LLM generator to generate future work (steps 6,7) (Task 3). Simultaneously, extracted future work from a paper is refined (re-extracted) by the LLM to remove sentences unrelated to future work (steps 1b,2b) and make ground truth (Task 2). Here, the ground truth of each paper contains a bunch of future work as a paragraph. Later, LLM-generated future work and ground truth are evaluated by the LLM (step 8, Figure 1), which assigns scores based on performance (step 9). If the generated output does not meet the threshold, the LLM re-generates it based on feedback (steps 10, 11) (Task 4).
Task 1: Section Selection. After extracting the texts (Section 3), we used a cosine similarity-based approach to identify the most relevant sections for generating future work. Each paper consists of sections such as Abstract, Introduction, and Methodology. We calculated the cosine similarity between each section and the paper’s future work section (Table 8, Appendix) across all papers, and we selected the top three sections with the highest average similarity, which were Abstract, Introduction, and Conclusion. These sections formed the basis for generating future work content.
Task 2: Refining Ground Truth. After extracting the future work sections described in Section 3, we employed an LLM as an extractor role to isolate only future work sentences while removing irrelevant sentences, resulting in a Refined future work (Figure 1, step 1b). Extracting the future work is challenging since most papers do not have a dedicated section, and it is often scattered in any other section. Therefore, filtering out noisy sentences is crucial to ensure accurate extraction ( 5.2). Additionally, we conducted a human evaluation to assess whether LLM is good at extracting future work sentences or not. By this process, we made silver standard data (Details in  6.4).
Task 3: Generating future work Using LLMs and RAG. Generating future work from a single paper may lead to narrow, redundant, or overly specific suggestions, missing broader research trends and interdisciplinary insights. Incorporating other papers’ content through RAG enhances contextual relevance, ensuring alignment with ongoing advancements and diverse, meaningful contributions reflective of real-world research priorities. To generate the future work of a paper, we utilized the top three sections (Abstract, Introduction, and Conclusion) selected in Task 1, integrating a RAG system, which consists of ‘X’ number of research papers (Details  5.3). The Retriever processes the input query, which includes the prompt and the top three sections of the research paper. It then retrieves additional relevant information from the vector database using cosine similarity and applies a ranking mechanism to prioritize the extracted text (Figure 1, steps 2a,3a,4,5). The final augmented input (a combination of the prompt, selected sections, and ranked retrieved text)is fed into the LLM Generator, producing future work content based on the provided context (Figure 1, step 6).
Task 4: Evaluation and Iterative Refinement. After generating future work from Task 3, we evaluated its quality against the Refined future work (ground truth from Task 2). We employed several NLP-based quantitative metrics and used an LLM as a Judge to assess LLM-generated text against the Refined future work based on coherence, relevance, readability, grammar, and overall quality (Figure 1, step 8,9). Each evaluation was rated on a 1 (worst) to 5 (best) scale, with justifications provided (Figure 4, Appendix). We set a threshold of 3 as an acceptable midpoint. If an LLM-generated future work received a score of 3 or lower in any metric, the justification was incorporated into the prompt, and the future work was regenerated accordingly. This iterative refinement process was repeated up to two times to assess whether performance improved (Figure 1, step 9,10 11).
Metrics w/o LLM’s feed w LLM’s feed
ROUGE-1 24.33 25.74(
+1.41
ROUGE-2 5.27 7.85(
+1.05
ROUGE-L 17.24 20.25(
+2.58
BScore(f1) 87.23 87.50(
+0.27
Jaccard S 15.40 18.13(
+2.73
Cosine S 48.07 57.03(
+8.96
BLEU 2.38 5.74(
+3.36
Coherence 3.94 3.97(
+0.03
Relevance 4.07 4.62 (
+0.55
Readability 3.19 3.49 (
+0.30
Grammar 4.02 4.01 (
-0.01
Overall 3.85 3.95 (
+0.10
Table 1: Comparison of performance without and with LLM feedback using GPT-4o mini. Metrics include Jaccard Similarity (Jaccard S) and Cosine Similarity (Cosine S).
Model It. Coh. Rel. Read. Gr.
Llama 3 ZS 1 3.83 4.47 4.04 3.44
Llama 3 FT 1 3.37 3.69 2.90 3.65
GPT 3.5 1 3.89 4.47 3.38 4.04
GPT 4om 1 3.94 4.07 3.19 4.02
GPT 4om + RAG 1 3.93 3.96 3.18 4.01
GPT 4om + RAG 2 3.97 4.50 3.50 4.06
GPT 4om + RAG 3 3.96 4.34 3.36 4.05
Table 2: Performance of different models in generating future work, evaluated against paper ground truth extracted by LLM (GPT Refined Ground Truth). Note: GPT 4om refers to GPT-4o mini, ZS indicates Zero-Shot, FT represents Fine-Tuning, It. denotes Iteration, Coh. stands for Coherence and Logic, Rel. for Relevance and Accuracy, Read. for Readability and Style, and Gr. for Grammatical Correctness.
Model Sec. Iter. R-1 R-2 R-L BS-F1 JS CS BLEU
RAW Ground Truth
BART 3 1 15.29 26.3 4.83 84.53 11.33 35.39 0.6
T5 3 1 - - - 17.84 1.85 5.86 0.0003
GPT 3.5 All 1 21.08 6.41 17.03 86.40 14.54 45.69 4.13
GPT 3.5 3 1 20.03 5.36 16.65 86.28 13.70 44.84 3.18
GPT 4o mini 3 1 17.50 3.11 14.22 85.91 11.55 40.53 1.29
GPT 4o mini + RAG 3 1 21.20 5.09 16.41 86.40 13.27 40.36 2.84
GPT Refined Ground Truth
GPT 3.5 3 1 28.17 8.68 20.64 87.88 18.49 55.36 5.61
GPT 3.5 3 2 27.80 8.67 20.49 87.88 18.66 56.72 5.81
GPT 4o mini 3 1 24.33 5.27 17.24 87.23 15.40 48.07 2.38
Llama 3 ZS 3 1 14.15 3.23 10.38 85.61 14.52 52.27 2.27
Llama 3 FT 3 1 26.75 5.61 - 85.54 12.28 45.37 1.14
GPT 4o mini 3 1 24.59 5.62 17.69 87.44 15.87 50.52 2.72
GPT 4o mini + RAG 3 1 21.13 5.52 14.19 87.67 16.95 49.32 3.42
GPT 4o mini + RAG 3 2 26.58 7.97 20.87 88.15 18.67 58.33 5.67
GPT 4o mini + RAG 3 3 26.96 7.53 20.18 87.94 17.92 56.34 4.68
Table 3: Performance comparison of various models in generating future work, considering Extracted future work from papers using Python regex (RAW Ground Truth, RAW GT) and LLM-refined future work extracted from RAW GT (GPT Refined Ground Truth, GPT GT). Note: Iter. denotes Iteration, Sec. refers to Section, R-1, R-2, and R-L represent ROUGE-1, ROUGE-2, and ROUGE-L scores, respectively. BS-F1 indicates BERTScore (F1), while JS and CS stand for Jaccard Similarity and Cosine Similarity, respectively.
5 Experimental Setup
5.1 Topic Modeling
For trend analysis, we employed topic modeling enhanced with an LLM to generate both topics and corresponding titles. Specifically, we integrated GPT 4o mini in a zero-shot configuration with BERTopic (since few-shot approaches tend to overfit by mirroring prompt examples) and set a minimum similarity threshold of 0.75.
5.2 Extraction
We employed the Science Parse tool and python regex to extract each paper’s future work sections. After isolating these texts, we applied GPT 4o mini with a temperature of 0 in a zero-shot configuration. Our prompt, “Extract sentences related to future work, nothing else,” successfully guided the model to extract a coherent paragraph comprised only of future work-related sentences, effectively discarding unrelated content.
5.3 Generation
To generate future work texts, we fine-tuned BART and T5 models using 70% training data and 30% testing data. BART has a maximum token limit of 1024, while T5 is limited to 512 tokens, and any input exceeding the respective token limits is disregarded. For both models, the generated text is capped at 100 tokens. Additionally, we applied GPT-3.5 and GPT-4o using a zero-shot approach with a set temperature of 0.2 for the same tasks. The prompts mentioned that the generated text shouldn’t be more than 100 words and leveraged their 16,000 and 128,000 token context windows, respectively.
Generating Cosine Similarity: We used Sentence Transformers (’all-MiniLM-L6-v2’) for embedding generation and scikit-learn’s  [^2] https://scikit-learn.org/stable/ cosine similarity function for similarity computation.
Fine-Tuning LLaMA: For LLaMA 3.1 fine-tuning, it was trained using formatted data, where ‘Abstract,’ ‘Introduction,’ and ‘Conclusion’ served as input, and extracted ‘future work’ was used as output. LLaMA already contains extensive general knowledge, so they require few domain-specific examples to adapt effectively. So, we split the dataset into 30% training and 70% testing. We applied LoRA Hu et al. (2022) with 4-bit quantization (QLoRA) Dettmers et al. (2023) for memory optimization, in LoRA, we set alpha=16 without dropout as default parameter. FlashAttention Dao et al. (2022) was incorporated to improve processing speed. We set a maximum input sequence length of 2048 tokens and trained the model for 60 steps with a default learning rate of 2e-4. During testing, we provided a prompt with detailed instructions for generating future work, with a maximum output length of 128 tokens and a temperature setting 1.
LLM with RAG Integration We integrated RAG with GPT-4o mini, leveraging OpenAI’s ‘text-embedding-3-small’ model. Our vector database comprises 100 randomly selected papers from the dataset, and these papers were removed from the dataset. For semantic search using vector embeddings, we employed LlamaIndex, utilizing its in-memory vector store rather than an external database. We segmented the data into chunks of up to 512 tokens to accommodate smaller context windows. The overall context window was set to 3,900 tokens—the maximum number of tokens the LLM can process at a time. Additionally, we set $K=3$, meaning that the top three most relevant chunks are retrieved from the vector store to provide contextual support during the generation process.
For LLM as a judge and LLM as an extractor, we used a one-shot approach with GPT 4o mini and a zero-shot approach with GPT 4o mini, respectively. We experimented with GPT 4o mini vs Llama 3-80b as a Judge and found Llama 3-80b provides a higher score than GPT 4o mini (11, Appendix).
6 Experiments and Results
6.1 Trend Analysis of future work through Topic Modeling
For trend analysis of future work, we chose BERTopic, and after testing various sentence transformer models, we selected ‘all-MiniLM-L6-v2’ (Table 16, Appendix). Additionally, we tested KMeans clustering with 15-30, but in all cases, silhouette and coherence scores dropped. We also experimented with UMAP, varying neighbors, and components (5–20), but its integration deteriorated performance (Figure 3, Appendix).
6.2 Evaluation of Extraction
Evaluation-raw ground truth vs. GPT-refined ground truth. We experimented with two ground truth approaches: (1) the raw ground truth (RAW GT) and (2) the GPT-refined ground truth (GPT GT) with LLM-generated text. RAW GT refers to the future work texts extracted from research papers using ScienceParse tool and regular expressions in Python. GPT GT takes RAW GT and utilizes the LLM to extract only sentences related to future work. After generating future work from key sections using our proposed approach, we evaluated ‘generated future work’ with RAW GT and GPT GT. As shown in (Table 6, Appendix), GPT GT outperformed RAW GT across all metrics, demonstrating improved text quality and relevance when using GPT 4o mini. Additionally, Table 3 confirms that GPT GT yields better performance than RAW GT across different models when comparing the same model’s RAW GT with GPT GT.
6.3 Evaluation of Generation
We experimented with various LLMs (Table 3) to generate future work sections. Traditional NLP-based metrics primarily focused on lexical overlap rather than semantic quality, while LLM-based evaluation provides a more contextual assessment than token-level comparison. Alongside the NLP-based evaluation, it focused on coherence and logic, relevance and accuracy, readability and style, grammatical correctness, and overall impression.
Evaluation-considering all sections vs top-3 sections. Instead of processing the entire paper, we used cosine similarity to select the three most relevant sections as input. As shown in (Table 4, Table 3), using only the top three sections resulted in a slight performance drop (e.g., ROUGE-L: -0.38, BERTScore: -0.12).
Metrics All sec 3 sec.
ROUGE-1 21.08 20.3(
-0.78
ROUGE-2 6.41 5.36(
-1.05
ROUGE-L 17.03 16.65(
-0.38
BScore(f1) 86.40 86.28(
-0.12
Jaccard S 14.54 13.70(
-0.84
Cosine S 45.69 44.84(
-0.85
BLEU 4.13 3.18(
-0.95
Table 4: Performance comparison between using three selected sections versus all sections as input for future work generation with GPT-3.5
Evaluation-incorporating LLM feedback. After generating the ‘future work’ text, a separate LLM served as a judge, scoring the output on a 1–5 scale with justification (feedback). This feedback was then integrated into the prompt to refine the generated text further. We evaluated the LLM-generated text after each iteration with GPT GT, and our results show a significant performance improvement across all metrics after incorporating LLM feedback once (Table 1). Moreover, from Table 3, we observed that after the first feedback loop (iteration 2), using GPT-4o mini + RAG, performance improved across nearly all metrics. However, in iteration 3, performance declined when the model received feedback a second time, indicating that excessive iterations introduced bias. These findings suggest that a single round of feedback integration is optimal for improving model performance without introducing unwanted biases.
Evaluation-among all models. We applied various models, and among all of the models, GPT 4o mini with RAG performs best when LLM feedback goes to the model (iteration 2) when considering ‘GPT Refined Ground Truth’ in NLP-based metrics (Table 3) and metrics in LLM as a Judge (Table 2). In terms of readability, Llama 3 zero shot performs better. We also compared our model (GPT 4o mini + RAG) with the simple zero shot model of GPT 4o mini, and we observed gains in Jaccard Similarity (+1.08), BLEU score (+0.7), and, most notably, BERTScore F1 (+0.23) (Table 3). However, n-gram-based metrics like ROUGE-L (-3.5) decreased due to the introduction of new information from RAG, which diluted exact text overlap.
6.4 Human Evaluation
Evaluating with only LLM-based evaluation metrics raises the question of reliability and potential biases. To overcome this problem, we compared LLM-based evaluation with human judgments and found a strong relationship. We conducted a user study with 200 randomly selected samples involving three annotators who rated below two questions from 1 to 3 (1 = worst, 3 = best). The annotators are graduate students, knowledgeable about machine learning, not related with this paper. The questions are: Q1: How well does LLM extract the future work (ground truth) without generating anything? Q2: How good has LLM generated future work (from Abstract, Introduction, and Conclusion) based on ground truth (paper mentioned future work)? Q3: How does the LLM Feedback approach improve performance in terms of originality? In Table 5, the average rating across all questions is approximately 2, reflecting an overall average performance. The rating for Q1 suggests that the LLM is effective at extracting text. Comparing LLM-generated future work with ground truth (Q2) shows average performance supported by strongly kappa and Kendall’s tau scores. Notably, in Q3, our proposed approach with one feedback iteration achieves an average human rating of 2.75, demonstrating that the LLM feedback loop significantly enhances result quality (Details A.1).
6.5 Trends in NLP Future Work
We analyzed the evolution of future work directions in NLP from 2012 to 2024 (Table 9, 7, Appendix). The early years emphasized neural machine translation, model improvement, and basic future work identification. From 2017 to 2019, there was a strong focus on language modeling, machine translation, and multilingual text representation, reflecting the rise of deep learning in NLP. From 2020 to 2021, research has expanded towards multi-scale learning techniques and generalization. In recent years, evaluation frameworks, bias mitigation, and performance robustness have increasingly been prioritized, highlighting the community’s shift toward more reliable, explainable, and fair NLP models. Early research focused on statistical methods, machine translation, and model improvement. Still, recent trends emphasize LLMs, prompt engineering, and data efficiency, also increasingly emphasize model performance robustness, privacy, adversarial attacks, generalization, and evaluation in NLP models (Figure 2, Table 14, 7 Appendix). From the topic analysis from the recent year (ACL 2024), we can see that the prominent topics are ‘linguistic,’ ‘LLM,’ ‘privacy,’ ‘annotation,’ etc. (Table 12, Appendix). Researchers also emphasize the need for better evaluation methods to assess and mitigate model biases, fairness, and societal implications of NLP models. Ensuring NLP models are trustworthy, explainable, and interpretable evaluation frameworks for document generation and annotation quality. Over time, the focus has shifted from a narrow concentration on model and translation improvements toward more nuanced themes that encompass ethical considerations, interdisciplinary approaches, and real-world applicability (Table 13, Appendix). In addition, Future research increasingly explore connections of interdisciplinary approaches between AI and human cognition.
Ques. Rating (Avg.) W. Kappa Kend’ Tau
Q1 2.54 0.24 0.29
Q2 1.92 0.30 0.36
Q3 2.75 0.28 0.29
Table 5: Average rating and annotators user agreement on user study.
7 Discussions
Our experiment shows that GPT-Refined Ground Truth aligns well with LLM-generated future work text, indicating that RAW Ground Truth contains more noise and irrelevant sentences. Our proposed approach can extract future work-related sentences from papers that lack explicit future work sections. Also, incorporating RAG with GPT 4o mini increased the Jaccard Similarity and BERTScore but reduced the n-gram overlaps matching than zero shot GPT 4o mini. Notably, one-time LLM feedback improves the performance of LLM when considering GPT Refined Ground Truth. To make generalizability, our model showed incorporating a single round of feedback led to significant improvements. However, applying a second round of feedback negatively impacted performance, producing a more biased response. We compared using all sections versus only the top three sections as input. Although restricting our analysis to the top three sections resulted in a minor performance drop (Table 4), we adopted this approach to mitigate the computational and API costs. The feedback might miss issues related to the originality of ideas or the soundness of the proposed future work. To alleviate this problem, we conducted a human experiment (Q3) and got a good average rating (2.75). We relied on the same LLM (GPT 4o mini) as the generator and evaluator, which enhanced the LLM’s own bias, we experimented with Llama 3 70b as an evaluator, but Llama provided unncessary higher scores even though sentence quality was not good (Table 11, Appendix).
8 Conclusions
The future work section is a forward-looking guide, helping the research community explore new directions. We selected the top three sections from each research paper to mitigate LLM context size limitations. We then utilized an LLM to extract future work, producing a more coherent ground truth that enhances model performance. Additionally, we integrated an external vector database to further improve LLM’s performance to generate future work from input text. For evaluation, we applied NLP-based metrics alongside an LLM-as-a-Judge approach, using explainable LLM metrics to assess performance, provide feedback, and iteratively refine text generation. Furthermore, we conducted a 13-year trend analysis to examine how future work priorities have evolved over time, highlighting key shifts in research focus.
Limitations and Future Work
The study focuses on ACL papers (2012–2024), providing strong domain relevance but limiting cross-disciplinary applicability. RAG retrieval was limited to 100 papers due to computational costs. Employing the same LLM as both generator and evaluator may introduce systematic biases that affect the output scores. Moreover, we assumed that the LLM-extracted Future Work could serve as the ground truth, that incorporating LLM feedback enhances performance, and that the suggestive Future Work generated by the LLM aligns with the ground truth—even though these evaluations were based on a limited number of annotators and samples. While the LLM evaluates coherence, reliability, readability, and grammatical correctness, these criteria may not fully capture essential academic qualities such as the originality of ideas or the robustness of the proposed future work and leads to overfitting. Furthermore, iterative refinement using the same model risks converging towards a narrow stylistic range and set of ideas, potentially increasing the likelihood of hallucinated content. For T5 and BART, we limited input length to 512 and 1024 tokens, respectively, which may restrict the model’s ability to capture complex research contexts fully. Also, we did not explore LLM feedback with Chain of Thought or self-consistency approaches, nor did we analyze hallucination levels with and without RAG integration. In the future, we will expand to broader research fields and refine extraction methods for implicit future work mentions. We will experiment on open-source LLMs to reduce API costs and conduct large-scale experiments to construct a comprehensive vector database to enhance retrieval quality in RAG. We will experiment with more human annotators to check whether the LLM as an extractor, generator, and iterative feedback loop sufficiently addresses deeper qualitative aspects of academic writing in the context of real-world applicability and make gold standard dataset. Moreover, to alleviate the bias from LLM feedback, we will experiment with Reinforcement Learning from Human Feedback (RLHF) and the AI Feedback (RLAIF) approach and make a more robust model. Also, we will investigate advanced LLM reasoning techniques, such as Chain of Thought and self-consistency, to refine LLM feedback mechanisms and expand future work extraction by incorporating cited papers and reviewers’ perspectives to diversify insights.
Ethics Statement
Our work involves extracting and generating future work sections from scientific articles, raising important ethical considerations regarding intellectual property, authorship, and responsible AI use. Our approach does not entirely rely on the automated generation of research papers; our framework uses the process to identify potential directions or gaps in the research. We also incorporated external evaluation metrics and human feedback to assess the quality of the generated content. It helps mitigate risks like overfitting to internal metrics or biases inherent in the LLM, ensuring that the final content is both robust and original. To ensure compliance with ethical research practices and academic integrity, we adhere to the following principles: 1. Respect for Original Authorship. We do not claim authorship of the generated future work. Instead, our method functions as an assistive tool for analyzing and refining research trends, complementing rather than replacing human intellectual contributions. 2. Fair Use and Transparency. The Extracted text is used only for research and analysis purposes. It is neither directly republished nor misrepresented, and we acknowledge the source articles when applicable to maintain full transparency. 3. Responsible AI Use & Avoiding Misleading Content. Our LLM-based generation process is designed to refine and organize existing content rather than fabricate entirely new research directions. We employ a RAG framework along with iterative refinement mechanisms to ensure that generated future work remains firmly anchored in the paper’s actual contributions. This strategy prevents speculative or arbitrary directions and mitigates potential guideline violations. In Section 6.4, we detail a user study involving three annotators who assessed the effectiveness of our approach in generating future work based on the paper’s mentioned content. Although the average rating of 1.92 on a scale from 1 (worst) to 3 (best) indicates moderate performance, it demonstrates that our method effectively grounds the output in validated, paper-specific information. 4. Non-Substitution of Human Contribution. The generated future work is intended to assist researchers by organizing, summarizing, and clarifying potential research directions. Final decisions and substantive content remain the responsibility of human authors, thereby preserving the integrity of scholarly work. 5. Alignment with Ethical Standards Our approach aligns with ACL’s ethical research guidelines and maintains rigorous academic integrity standards. We do not automate the entire writing process; instead, we offer a tool to help researchers better structure and refine their future work suggestions.
Appendix A Appendix
Metrics RAW GT GPT GT
ROUGE-1 17.50 24.59(
+7.09
ROUGE-2 3.11 5.62(
+2.51
ROUGE-L 14.22 17.69(
+3.47
BScore(f1) 85.91 87.44(
+1.53
Jaccard S 11.55 15.87(
+4.32
Cosine S 40.53 50.52(
+9.99
BLEU 1.29 2.72(
+1.43
Table 6: Comparison of performance between extracted ground truth from papers (RAW GT) and LLM-refined ground truth (GPT GT) using GPT-4o mini.
Topics Year
Model Improvement (31), Diffusion (13) 2012
Image Captioning (54), Machine Translation (syntax) (12) 2013
Sentiment Analysis (11) 2014
Machine Translation (20), Opinion Mining (10) 2015
NLP models (85) 2016
Multilingual Text (110) 2017
Neural Machine Translation (216) 2018
NLP Models (265) 2019
Neural Machine Translation (49), Bias in Language Models (LM) (21), Document Generation (21), Conversational QA (19), Dialogue Evaluation (16), Adversarial Attacks and (13), Named Entity Recognition (12) 2020
Generalization and Robustness (36), Multilingual Neural Machine Translation (33), Robustness (28), Conversational AI Evaluation (26) 2021
Machine Translation (63), Model Training and Evaluation (62), Multidisciplinary (30), NLP and Event Detection (28), Multilingual (27), Knowledge Augmentation in LM (24), Annotation Quality (19), Summarization in Educational Content (14), Funding (13) 2023
LM Performance Evaluation (147), Performance and Data Evaluation (61), Advancing LLMs (66), Logic and Relationships (41), Model Performance, Privacy and Model Defense (28), Common Sense Reasoning and Evaluation Methods (19), Funding (11) 2024
Table 7: Prominent Topics of Each Year (2012-2024). Note: Number means how many time these topics occurred.
Figure 2: Evolving future work Over Years (2012-2024)
Model Abs. Intro. RW Data Meth. Exp. Con. Lim.
Cos. Sim 25.08 24.52 21.09 8.96 14.82 21.79 22.0 21.38
Table 8: Average Cosine Similarity of each section with future work. (Abs., Intro., RW, Data, Meth., Exp., Con., and Lim. indicate Abstract, Introduction, Related Work, Dataset, Methodology, Experiment and Results, Conclusion, and Limitation
Dataset future work
2012 Future Research in Natural Language Processing and Model Improvement Techniques
2013 future work in Natural Language Processing and Image Captioning
2014 Future Directions in NLP Model Development and Sentiment Analysis
2015 Sentiment Analysis and Dispute Detection Models in Conversational Text
2016 Future Directions in Language Modeling and Machine Translation
2017 Future Research Directions in NLP Models and Techniques
2018 Future Directions in Model Development for Multilingual Text Representation
2019 Neural Machine Translation and Future Research Directions in NLP Models
2020 Future Research Directions in NLP Models and Methods
2021 Future Research Directions in Neural Machine Translation and Multi-Scale Learning Techniques
2022 Evaluation Frameworks and future work in Generalization and Robustness
2023 Future Directions in NLP Model Evaluation and Bias
2024 Future Directions in Language Model Research and Performance Evaluation
Table 9: Prominent future work in Each year.
A.1 Human Evaluation
We collected 200 random samples from the dataset and, for each sample, evaluated three key questions using Ground Truth, LLM-generated text, and LLM-regenerated text after feedback. Table 10 in the Appendix shows that each question is divided into two columns (A and B) representing different text types. Specifically, in Q1, human annotators compared the future work extracted directly from the paper (Ground Truth) using tools with the future work obtained via LLM extraction, assessing the LLM’s ability to accurately extract content without generating new text. In Q2, the evaluation focused on how well the LLM generated future work from input texts (Abstract, Introduction, and Conclusion), with the paper mentioning future work (ground truth). Finally, Q3 examines the impact of incorporating an iterative feedback loop by comparing the output of the LLM + RAG method with and without the additional feedback iteration. This comparison highlights how the feedback mechanism contributes to refining and improving the final output.
Quest Column A Column B
Q1 Ground Truth (Extracted text from paper using parsing tool) Ground Truth (Extracted text using LLM)
Q2 Ground Truth future work LLM Generated future work
Q3 Generated future work by LLM + RAG Generated future work by LLM + RAG with Feedback
Table 10: Human Evaluation Details
Model Coherene Relevance Readability Overall Impression
GPT 4o mini 3.89 4.47 3.38 3.82
Llama 3-80b 4.62 4.37 4.81 4.49
Table 11: Performance between considering GPT 4o mini vs Llama-3 80b as a Judge
future work Title
euphemisms, euphemism’, turkish, linguistic, multilingual
aspectual, llm, feature, llms, hyperparameters
benchmarks, pipeline, hyperparameters, modeloutput, accuracy
future, goals, satisfaction, achievement, achieve
privacy, datasets, annotation, data, identifiable
semantic, commonsense, semantics, annotation, schema
ai, mobility, future, intelligence, nrrp
Table 12: Prominent future work topics of ACL 2024 papers.
future work Title Count
2024
Future Directions in Language Model Research and Performance Evaluation 147
Future Research Directions for Advancing LLMs and Prompt Engineering 66
Future Research Directions in Model Performance and Data Efficiency 61
Future Research on Logic and Relationships 41
Future Directions in Privacy and Model Defense Research 29
Future Directions in Commonsense Reasoning and Evaluation Methods 19
Project Funding and Support for AI Research 11
2023
Simultaneous Machine Translation Research 63
Future Directions in Model Training and Evaluation 62
Future Research Directions in Multidisciplinary Approaches 30
Future Directions in NLP and Event Detection 28
Future Directions in Multilingual Research and Model Evaluation 27
Future Directions in Knowledge Augmentation for Language Models 24
Annotation Quality and Future Improvements in Predictive Tasks 19
Future Directions in Lay Summarization and Educational Content Analysis 14
AI Research Funding and Support Initiatives 13
2021
Evaluation Frameworks and future work in Generalization and Robustness 36
Future Directions in NLP Research and Applications 35
Future Directions in Multilingual Neural Machine Translation Research 33
Robustness and Learnability in NLP Models 28
Future Directions in Conversational AI and Evaluation Systems 26
Future Multimodal Research in Visual Tasks and Event Predictions 19
2020
Future Research Directions in Neural Machine Translation and Multi-Scale Learning Techniques 49
Lack of future work Mentioned in Provided Texts 26
Future Research on Language Models and Biases in Text Analysis 21
Future Research Directions in Document Generation and Understanding 21
Knowledge Base Question Answering and Future Directions in Conversational QA Systems 19
Future Directions in Open-Domain Dialogue Evaluation and Model Development 16
NLP Vulnerability and Robustness in Adversarial Attacks and Counterspeech Generation 13
Discontinuous Named Entity Recognition and Ontology Alignment in Information Extraction 12
2019
Future Research Directions in NLP Models and Methods 265
future work Identification in Texts 27
2018
Neural Machine Translation and Future Research Directions in NLP Models 216
future work Identification and Analysis 29
2017
Future Directions in Model Development for Multilingual Text Representation 110
future work Identification and Analysis 15
2016
Future Research Directions in NLP Models and Techniques 85
Lack of Identified future work in Provided Texts 13
2015
Future Directions in Language Modeling and Machine Translation 20
Opinion Mining and Relation Modeling in NLP 10
2014
Future Directions in NLP Model Development and Sentiment Analysis 11
2013
future work in Natural Language Processing and Image Captioning 54
future work in Syntax-Based Statistical Machine Translation Improvements 12
2012
Future Research in Natural Language Processing and Model Improvement 31
future work in Diffusion Prediction 13
Table 13: Overview of recurring future work topics in ACL papers for each year, showing the number of times each topic appears. The count represents the frequency of mentions across various research papers.
Hierarchical Future Directions Category
NLP Model Development 1
Improve models and Techniques
Sentiment Analysis
Language Modeling and Machine Translation
Model Development for Multilingual Text Representation
Improve Neural Machine Translations (NMT)
NMT with Multilingual Text Representation
NMT with Multi-scale Learning Techniques
Model Tuning with Multimodal Performance Analysis
Language Model Research with Performance Evaluation
Model Tuning with Multimodal Performance Analysis
Language Model Research and Performance Evaluation
Advancing LLMs and Prompt Engineering
Model Performance and Data Efficiency
Machine Translation and Multilingual Models 2
Syntax-Based Statistical Machine Translation Improvements
Sentiment Analysis and Relation Modeling 3
Sentiment Analysis and Dispute Detection Models in Conversational Text
Opinion Mining and Relation Modeling in NLP
Question Answering (QA) and Conversational Systems 4
Knowledge Base Question Answering and Future Directions in Conversational QA Systems
Future Directions in Open-Domain Dialogue Evaluation and Model Development
Conversational AI and Evaluation Systems
Retrieval-Based QA and Summarization
Robustness and Vulnerability in NLP 5
NLP Vulnerability and Robustness in Adversarial Attacks and Counterspeech Generation
Evaluation Frameworks and future work in Generalization and Robustness
Robustness and Learnability in NLP Models
Future Directions in Privacy and Model Defense Research
Information Extraction and Analysis 6
Discontinuous Named Entity Recognition and Ontology Alignment in Information Extraction
Multimodal and Visual Tasks 7
future work in Natural Language Processing and Image Captioning
Future Multimodal Research in Visual Tasks and Event Predictions
Future Research Directions in Multimodal Performance Analysis
Bias, Ethics, and Societal Impact 8
Future Research on Language Models and Biases in Text Analysis
Future Directions in NLP Model Evaluation and Bias
Document Generation and Annotation 9
Future Research Directions in Document Generation and Understanding
future work in Annotation Quality and Evaluation Techniques
Psychology and AI 10
Future Research Directions and Methodological Extensions in AI and Psychology
Logic and Relationships in AI 11
Future Research on Logic and Relationships
Evaluation Methods and Metrics 12
Future Directions in Commonsense Reasoning and Evaluation Methods
Support and Funding for AI Research 13
Project Funding and Support for AI Research
Table 14: Hierarchical future work.
Figure 3: BERTopic + LLM for generating title in future work.
Instructions: You are provided with two texts for each pair: one is generated by a machine (Machine-Generated Text), and the other is the original or ground truth text (Ground Truth). Please read both texts carefully. After reviewing each text, assign a score from 1 to 5 based on the criteria outlined below. The score should reflect how well the machine-generated text compares to the ground truth, where 1 represents poor quality and 5 represents excellent quality that closely matches or even surpasses the ground truth in some aspects. Scoring Criteria: Coherence and Logic: 5: The text is exceptionally coherent; the ideas flow logically and are well connected. 3: The text is coherent but may have occasional lapses in logic or flow. 1: The text is disjointed or frequently illogical. Relevance and Accuracy: 5: The text is completely relevant to the topic and accurate in all presented facts. 3: The text is generally relevant with minor factual errors or slight deviations from the topic. 1: The text often strays off topic or includes multiple factual inaccuracies. Readability and Style: 5: The text is engaging, well-written, and stylistically consistent with the ground truth. 3: The text is readable but may lack flair or have minor stylistic inconsistencies. 1: The text is difficult to read or stylistically poor. Grammatical Correctness: 5: The text is free from grammatical errors. 3: The text has occasional grammatical errors that do not impede understanding. 1: The text has frequent grammatical errors that hinder comprehension. Overall Impression: 5: The text is of a quality that you would expect from a professional writer. 3: The text is acceptable but would benefit from further editing. 1: The text is of a quality that needs significant revision to be usable. Task: For each text pair: Rate the Machine-Generated Text on each criterion and provide a final overall score out of 5. Provide a brief justification for your scores, highlighting strengths and weaknesses observed in the machine-generated text relative to the ground truth. '''   Example of Usage: Text Pair 1: Machine-Generated Text: "The quick brown fox jumps over the lazy dog repeatedly." Ground Truth: "A quick brown fox consistently jumps over the lazy dog." Evaluation: Coherence and Logic: 5 Relevance and Accuracy: 4 Readability and Style: 5 Grammatical Correctness: 5 Overall Impression: 5 Justification: The machine-generated text maintains the core message and style of the ground truth, presenting it coherently and engagingly. Minor variations in wording do not impact the overall quality or relevance of the message. """
Figure 4: Prompt for evaluation (LLM as a Judge).
Your task is to generate a refined "future work" section for a scientific article. Below, you will find sections from a scientific article including the ’Abstract’, ’Introduction’, ’Conclusion’ of a scientific paper. The goal is to ensure that this section clearly outlines future research directions without the issues highlighted below. Please focus on maintaining coherence, relevance, readability, grammatical correctness, and overall quality, ensuring the text flows logically and stays directly aligned with the main topic. I am providing the texts and found these problems. At first read the feedback and try to improve them when you generte future work. [LLM Feedback] Based on these details, please generate comprehensive and plausible future work suggestions that could extend the research findings, address limitations, and propose new avenues for exploration. future work should be within 100 words.
Figure 5: Incorporating LLM Feedback to the Model.
Dataset Coherence score Silhouette score
ACL 12 0.63 0.66
ACL 13 0.58 0.38
ACL 14 0.61 0.46
ACL 15 0.63 0.45
ACL 16 0.55 0.47
ACL 17 0.54 0.56
ACL 18 0.53 0.49
ACL 19 0.53 0.64
ACL 20 0.52 0.73
ACL 21 0.57 0.62
ACL 22 0.60 0.53
ACL 23 0.64 0.51
ACL 24 0.59 0.38
Table 15: Performance of each year’s future work.
Model Silhouette score
allenai-specter 0.558131
paraphrase-MiniLM-L6-v2 0.801949
all-mpnet-base-v2 0.732606
all-MiniLM-L6-v2 0.819191
all-MiniLM-L12-v2 0.479191
paraphrase-multilingual-MiniLM-L12-v2 0.815246
paraphrase-MiniLM-L6-v2 0.803524
bert-base-nli-mean-tokens 0.806526
all-distilroberta-v1 0.488876
msmarco-distilbert-dot-v5 0.436424
multi-qa-MiniLM-L6-cos-v1 0.568171
Table 16: Performance of various Sentence Transformers
LLM extracted Ground Truth future work LLM Generated future work. BS (p) BS (R) BS (f1)
future work should also include exploring the impact of different hyperparameters on model performance and investigating the potential benefits of incorporating external knowledge sources into the model. Furthermore, studying the generalizability of the model to different languages and domains would be an important direction for future research. Additionally, investigating the interpretability of the model’s predictions and exploring ways to improve model explainability could also be valuable avenues for further study." future work could focus on exploring novel architectures that combine the strengths of both encoder-only and decoder-only language models to improve semantic understanding in natural language processing tasks. Additionally, research could investigate alternative training methods or prompting techniques to enhance the capabilities of decoder-only models in comprehending word meaning. Furthermore, expanding the study to include languages other than English and evaluating a wider range of language models could provide a more comprehensive understanding of the performance differences between encoder-only and decoder-only architectures. Investigating the impact of model size and training data on semantic understanding could also be a valuable direction for future research. 0.87 0.90 0.88
Table 17: LLM Extracted vs LLM Generated future work (Note: BS, p, R, f1 indicate BERTScore, Precision, Recall, and F1 score.